A poster must have 32 square inches of printed matter with margins of 4 inches at the top and bottom, and 2 inches at each side. Find the dimensions of the whole poster if its area is minimum.



1
Expert's answer
2021-01-14T17:38:35-0500

"\text{let }x,y\text{ length and width of the poster}"

"S = xy"

"S_{prn}=(x-8)(y-4)=32"

"y= 4 +\frac{32}{x-8}"

"S(x)= 4x +\frac{32x}{x-8}"

"\text{find the extremum points where the derivative is 0}"

"S^{'}(x)= 4 +\frac{32(x-8)-32x}{(x-8)^2}=4-\frac{256}{(x-8)^2}"

"4-\frac{256}{(x-8)^2}=0"

"4=\frac{256}{(x-8)^2}"

"(x-8)^2 =64"

"\text{as } x>0 \ x=16"

"S^{'}(10)<0;S^{'}(20)>0"

"\text{then }x=16 \text{ minimum point}"

"y= 4+\frac{32}{16-8}=8"

Answer:16 inches and 8 inches poster sizes